13. Write the slope-intercept equation for the line that passes through (-5, -8) and is perpendicular to 10x – 6y = -11.
Answer by mananth(16949)

10 x -6 y = -11
Find the slope of this line
-6 y = -10 x -11
Divide by -6
y = 1 2/ 3 x + 1 5/6
Compare this equation with y=mx+b
slope m = 5/3

The slope of a line perpendicular to the above line will be the negative reciprocal -3/5
m1*m2=-1


m= - 3/5 ,point ( -5 , -8 )
Find b by plugging the values of m & the point in
y=mx+b
-8 = 3.00 + b
b= -11
m= - 3/5
The required equation is y = - 3/ 5 x -11
m.ananth@hotmail.ca